OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

@zepdrix

OpenStudy (dayakar):

equation of the line that is parallel to the line 4x - 3y = -12 is in the form of 4x-3y = k ,(k is a constant) to find k value substitute and the point (-3, 4).in the above equation,we get \[4\times \left( -3 \right)-3\times4=k\] -12-12=k -24 = k therfore ,k= 24 Required equation is 4x-3y=24